OpenStudy (anonymous):

Mary has a bag that contains 5 green marbles, 7 red marbles and 9 blue marbles. Without looking, Mary draws a marble from the bag. What is the probability that Mary will draw a red or a green marble? a.4/7 b.1/20 c.5/63 d.1/12

OpenStudy (irishboy123):

Red + green = 12 Total = 21 Got it?!

OpenStudy (anonymous):

5 + 7 + 9 = 21. \[\frac{ 5 }{ 21 } \times \frac{ 7 }{ 21 }\] reduce if needed. \[\frac{ 5 }{ 21 }\times \frac{ 1 }{ 3 }\] Solve and you have your answer. :)

OpenStudy (anonymous):

That is if the problem is asking P(red, then green). :)

OpenStudy (anonymous):

5/63

OpenStudy (anonymous):

Correct! :D

OpenStudy (anonymous):

if g(x)=x^2+3x-7 find g(-3) a. 11 c. –43 b. 7 d. –7

OpenStudy (anonymous):

@Friends1

OpenStudy (anonymous):

Hmmm....hold on. Haven't done these type of problems in quite some time. :)

OpenStudy (anonymous):

ok

OpenStudy (zarkon):

you might want to look at that probability question again